OpenStudy (anonymous):

How much greater is the surface area of a cube when the length is multiplied by 5

OpenStudy (anonymous):

A=6*x*x=6x^2 A'=6*(5x)*(5x)=25*6*x*x=25*6x^2 new one 25 time greater than first area

OpenStudy (anonymous):

Thank you :)